OpenStudy (anonymous):

multiply....(x-9/4)(x+4/9)

OpenStudy (vishweshshrimali5):

\[\frac{(x-9)}{4}\frac{(x+4)}{9}\] \[\frac{(x-9)(x+4)}{36}\] \[\frac{(x^2-9x+4x-36)}{36}\] \[\frac{x^2-5x-36}{36}\]